Q-2): Find the solution of the differential equation: y – 10y + 25 = 2 e5x + sin(x)
Answer #1

solution The given diffenertial y-10y + 25 = 2 **+ Son ( equation is (i) ma be a Let y = e(where being Constant) solution ofIntegral of Now the particular -(i) is (26 5 + sinta Dr _ IOD + 25 (205) D-LOD+25 sin(n) D²-LOD+25 5n 2e - [0*100+25] (205)50 = 2e 5* [02 -/ [0--200 +100] (eo) **00** 2 5) - 100}{0*+25)100 [(D*+25) +100] (sin() - 2 e Les | 0-20407100) [co-25) = 100е 52 -1 50 + [ 676 (or (sinn) + 25 sinn flod op (sinx) 57 10 Sina +25 sina + lo cosa ul + [- 50 676 - e50 + 50 24 sina flocos